OpenStudy (zephyr141):

Electromagnetism: Point-like charges Can I get some help on how to start this. My notes aren't really helping me right now.

OpenStudy (jiteshmeghwal9):

What do u want to learn?

OpenStudy (zephyr141):

im trying to find the electric field at an arbitrary point.

OpenStudy (zephyr141):

q1=3uC, q2=2uC and q3=-3uC. they're located at R1=(0 0 0), R2=(0 3cm 0) and R3=(0 0 2cm) respectively. find the electric field at an arbitrary point R=(x y z) using the general result, calculate the field at the point (5cm, 5cm, 5cm)

OpenStudy (zephyr141):

so i guess step one i'll have to convert cm to m so that's R2=(0, 0.03m, 0) and R3=(0,0,0.02m) and the position of (0.05m, 0.05m, 0.05m)

OpenStudy (zephyr141):

gonna use this as well\[F=k_e \frac{ q_1q_2 }{ r^2 }\]

OpenStudy (jiteshmeghwal9):

The arbitrary point 'R' is (x,y,z). Distance of point 'R' from charge \(q_1\)=\(\sqrt{x^2+y^2+z^2}\) Distance of point 'R' from charge \(q_2\)=\(\sqrt{x^2+(y-0.03)^2+z^2}\) Distance of point 'R' from charge \(q_3\)=\(\sqrt{x^2+y^2+(z-0.02)^2}\)

OpenStudy (jiteshmeghwal9):

I guess u r having problem in finding the resultant of the electric fields.

OpenStudy (zephyr141):

i think so. I need to refresh my memory on vectors and such. it's been a year since my last physics class.

OpenStudy (zephyr141):

this is the page from my notes. just trying to make sense of it.

OpenStudy (jiteshmeghwal9):

@Irishboy123 he may help u

OpenStudy (jiteshmeghwal9):

actually i learnt 3-d geometry 6 months ago

OpenStudy (zephyr141):

thank you for your help so far.

OpenStudy (jiteshmeghwal9):

:)

OpenStudy (zephyr141):

so for the force of q1 on q2 i get 59.93N and for q1 to q3 i'm getting -202.27N but i think that may be wrong... or all of this is wrong....

OpenStudy (zephyr141):

and the force of q3 on q2 is -41.62N

OpenStudy (ljetibo):

The resulting field will just be the individual forces from each charged particle. Since @jiteshmeghwal9 was so kind to do the distances for us the answer is just simply: $$E_{tot} = E_1 + E_2 + E_3 = \frac{1}{4\pi\epsilon_0}\left[ \frac{q_1}{\sqrt{x^2+y^2+z^2}} + \frac{q_2}{\sqrt{x^2+(y-0.03)^2+z^2}}+ \\ \frac{q_3}{\sqrt{x^2+y^2+(z-0.02)^2}} \right]$$ To get the field in (5, 5, 5) all you have to do is put in x=5, y=5, z=5. Because this is expressed in m then convert the 5cm to m too.

OpenStudy (jiteshmeghwal9):

Electric field is a vector quantity we cannot use simple algebraic methods to get the resultant

OpenStudy (zephyr141):

so it's the last part of my notes where it uses the sigma.

OpenStudy (ljetibo):

You can if you're only interested in the magnitude of the electric field as was implied by some of the answers of OP, but you are right - not in general. I can do the vector calcs a bit later.

OpenStudy (zephyr141):

so it's saying magnitude AND direction. so i guess we'll need to do some calculus. in my notes i think i see what i'm supposed to use.

OpenStudy (zephyr141):

this picture is a little closer but right there down at the bottom, i think that's what i'm supposed to use to solve this problem. do you agree?

OpenStudy (zephyr141):

or rather in the ovals.

OpenStudy (zephyr141):

ok nevermind i see what's up now.

OpenStudy (ljetibo):

that is what you need to do yes. The indices ij don't really "mean" all that much they just indicate that you have to do the sum by all pairs of charges and skip the rii situations (particle acting on itself, to avoid dividing by zero) because the distance then is zero.